11.Two fair six sided dice are rolled (i) What is the probability that the smallest of the two results is 3 given that the sum (ii) What is the probability that the sum of the two results is at most 5 given that the ii) What is the probability that the sum of the two results is 7 given that exactly one of the two results is 8? number 2 appeared at least once? of the two results is odd?I want to see the thought process behind this so that I could approach similar problems on my own, so please attempt to show every step and formula used, thank you.
Answer #1

a)

P(smallest is 3 | sum is 8)

P(A |B) = P(A and B)/P(B)

now

A = smallest is 3 and B is sum is 8

event A and B is

(2,6),(3,5),(4,4),(5,3),(6,2)

total number = 5

3 is smallest in 2 cases   {(3,5),(5,3)}

hence

P(smallest is 3 | sum is 8)

= P(smallest is 3 and sum is 8)/ P(Sum is 8)

= 2/5

b)

P(sum <= 5 | 2 has appeared at least once)

= P(sum <= 5 and 2 has appeared at least once)/P(2 has appeared at least once)

there are 11 possibilities when 2 has appeared at least once

sum <= 5 and 2 is there at least once

(1,2),(2,1),(2,2),(2,3),(3,2)

total 5 cases

hence

probability = 5/11

c)

sum is 7 | exactly one is odd

sum is 7 and exactly one is odd =

(1,6),(2,5),(3,4),(4,3),(5,2),(6,1)

total 6 cases

exactly one is odd = 18 cases

hence

6/18

=1/3

Please rate